# SciKnowEval: Evaluating Multi-level Scientific Knowledge of Large Language Models

### v3

@@ Abstract
Large language models (LLMs) have gained increasing prominence in scientific research, but there is a lack of comprehensive benchmarks to fully evaluate their proficiency in understanding and mastering scientific knowledge. To address this need, we introduce the SciKnowEval benchmark, a novel framework that systematically evaluates LLMs across five progressive levels of scientific knowledge: studying extensively, inquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. These levels aim to assess the breadth and depth of scientific knowledge in LLMs, including memory, comprehension, reasoning, discernment, and application. Specifically, we first construct a large-scale evaluation dataset encompassing 70K multi-level scientific problems and solutions in the domains of biology, chemistry, physics, and materials science. By leveraging this dataset, we benchmark 26 advanced open-source and proprietary LLMs using zero-shot and few-shot prompting strategies. The results reveal that despite the state-of-the-art performance of proprietary LLMs, there is still significant room for improvement, particularly in addressing scientific reasoning and applications. We anticipate that SciKnowEval will establish a standard for benchmarking LLMs in science research and promote the development of stronger scientific LLMs.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated an impressive capability in storing and recalling world knowledge, continuously expanding the boundaries of artificial intelligence. Their exceptional performance has permeated diverse specialized domains, including the scientific domain, leading to the emergence of scientific LLMs, such as Galactica [35], SciGLM [45], and ChemLLM [46]. To steadily advance scientific research, it is crucial to establish reliable benchmarks that comprehensively evaluate these models’ capability in handling scientific knowledge.
@@
While several existing LLM benchmarks [23, 48, 5] have incorporated scientific questions into their evaluations, and some benchmarks [34, 38, 4, 39, 25, 14] are specifically tailored for the scientific domain, we argue that the current benchmarks do not fully evaluate the potential of LLMs in scientific research due to their inherent limitations. Firstly, many existing benchmarks, such as AGIEval [48], SciQ [39], and ScienceQA [25], include science questions only up to the high school level, failing to tap into the deeper capability of LLMs. Secondly, recent scientific domain benchmarks like ChemLLMBench [14], SciBench [38], and SciAssess [4], despite involving more specialized scientific tasks, lack a comprehensive evaluation system, resulting in a limited understanding of capabilities. Lastly, most benchmarks overlook the assessment of safety issues in scientific research, even those attempting a multi-dimensional comprehensive evaluation such as SciEval [34].
In response to these deficiencies, in this study, we adopt a distinctive perspective, “LLMs as Scientists”, to revisit the evaluation in the scientific domain. We draw inspiration from the profound principles of Confucius outlined in the ancient Chinese philosophy “Doctrine of the Mean” , and present a novel Scientific Knowledge Evaluation benchmark, referred to as SciKnowEval, as illustrated in Fig. 1. This benchmark aims to assess LLMs based on their proficiency in five progressive levels: studying extensively, enquiring earnestly, thinking profoundly, discerning clearly, and practicing assiduously. Each level offers a unique perspective on evaluating the capabilities of LLMs in handling scientific knowledge, including memory, comprehension, reasoning, discernment, and application.
@@
Different from the widely adopted Bloom’s Taxonomy framework [20], we aim to encourage a balanced and comprehensive view of knowledge and its application, fostering a model’s capacity to think and act responsibly. In comparison to existing benchmarks, SciKnowEval mainly has the following characteristics: (1) It designs a systematic scientific knowledge evaluation framework that encompasses five progressive levels to mirror the learning process of humans. (2) It uses data from diverse sources, including scientific textbooks, literature, and databases, making it diverse and large-scale. (3) It places significant emphasis on scientific ethics and safety while comprehensively evaluating capabilities. Table 1 shows the detailed comparison of SciKnowEval with other benchmarks.
Figure 1: Illustration of SciKnowEval. We consider LLMs as scientists and evaluate their performance based on a hierarchical framework inspired by the ancient Chinese philosophy from “ Doctrine of the Mean”, encompassing five progressive levels that span from knowledge memory to application. We also construct a large-scale dataset comprising various scientific domain tasks at different levels, along with corresponding evaluation criteria, leading to a reliable benchmark of LLMs.
@@
SciKnowEval represents a comprehensive benchmark for assessing the capability of LLMs in processing and utilizing scientific knowledge. It aims to promote the development of scientific LLMs that not only possess extensive knowledge but also demonstrate ethical discernment and practical applicability, ultimately contributing to the advancement of scientific research. We summarize the contributions of this paper as follows:
@@
  • We construct a large-scale evaluation dataset comprised of 70K diverse scientific problems from the domains of biology, chemistry, physics and material science, accompanied by corresponding solutions and evaluation metrics, facilitating an extensive assessment of the breadth and depth of scientific knowledge encapsulated in LLMs.
  • We evaluate a wide range of advanced LLMs (including 18 general-purpose LLMs and 8 scientific LLMs) and rank their performance with the SciKnowEval dataset, elucidating both their strength and weaknesses.
Table 1: Comparison between SciKnowEval and other benchmarks.
[…]
L1 L2 L3 L4 L5
MMLU_sci STEM ✓ ✓ MCQ Exam, Book, Course 14,042
AGIEval_sci Phys., Geo., Bio., Chem., Math ✓ ✓ MCQ Exam 8,062
ARC_sci STEM ✓ ✓ MCQ Exam 3,548
SciQ Phys., Bio., Chem. ✓ ✓ MCQ ✓ Crowdsourcing 1,000
ScienceQA Natural Science ✓ ✓ MCQ, T/F Online Platform 4,241
SciBench Phys., Chem., Math ✓ GEN ✓ Textbook 789
SciEval Phys., Bio., Chem. ✓ ✓ ✓ MCQ, T/F GEN Website, Knowledge base 15,901
SciAssess Chem., Bio., Drug, Mat. ✓ ✓ MCQ, T/F RE, GEN Literature, Knowledge base 1,579
SciMT-Safety Chem., Bio., Drug ✓ GEN ✓ User queries 432
ChemBench Chem. ✓ ✓ ✓ ✓ MCQ, GEN ✓ Exam, Book Knowledge base 7059
LAB-Bench Bio. ✓ ✓ ✓ MCQ ✓ Literature, Knowledge base 2400
SciKnowEval Bio., Chem., Phys., Mat. ✓ ✓ ✓ ✓ ✓ MCQ, T/F RE, GEN Literature, Textbook, Database 70,203
2 Related Works
Assessing proficiency in scientific knowledge is a crucial aspect of LLM evaluation. For example, renowned benchmarks like MMLU [16], AGIEval [48], and ARC [5] have incorporated a number of scientific questions into their assessments. Recently, with the rapid application of LLMs in science research, several benchmarks tailored for scientific domains have been developed. SciQ [39] encompasses 13K crowdsourced science examination questions, covering diverse subjects including Physics, Chemistry, Biology, and more. ScienceQA [25] consists of 21K multimodal scientific questions and answers (QA) collected from elementary and high school science curricula, focusing on evaluating the interpretability of LLMs in addressing scientific problems. SciBench [38] collects open-ended questions from college-level textbooks in physics, chemistry, and mathematics to assess reasoning abilities for complex scientific problems. SciEval [34], based on Bloom’s Taxonomy, offers a multidisciplinary, multi-level dataset to assess understanding, application, computation, and research abilities. SciAssess [4] introduces a multimodal scientific literature analysis dataset, encompassing tasks like information extraction and chart/table QA. SciMT-Safety [15] presents a red-teaming benchmark with 432 malicious queries, focusing on potential misuse risks of LLMs. ChemBench [27] establishes a comprehensive benchmark consisting of 7,059 questions designed to assess the expertise and safety of LLMs across 11 subfields of chemistry. LAB-Bench [22] introduces the language agent biology benchmark, which includes over 2,400 multiple-choice questions aimed at evaluating AI systems on a range of practical biology research capabilities.
Table 1 has reported the statistics of these benchmarks. One can observe that they suffer from several limitations, such as inadequate scope of ability examination, limited data size, and insufficient task diversity. In this study, we develop a systematic scientific knowledge evaluation framework and construct a large-scale dataset with multi-level domain tasks to benchmark the general and scientific LLMs.
3 The SciKnowEval Dataset
@@ 3.1 Design Philosophy
The design philosophy of SciKnowEval is inspired by the profound principles of Confucius elucidated in the ancient Chinese book “Doctrine of the Mean”: Studying extensively, Enquiring earnestly, Thinking profoundly, Discerning clearly, and Practicing assiduously. This principle reflects the five progressive levels in the human learning process. In this study, we regard LLMs as Scientists and utilize this concept to evaluate them. Specifically, each level provides a perspective to assess the proficiency of LLMs, as described below.
  • L1: Studying extensively (i.e., knowledge memory). This dimension evaluates the breadth of an LLM’s knowledge across various scientific domains. It measures the model’s ability to remember a wide range of scientific concepts.
  • L2: Enquiring earnestly (i.e., knowledge comprehension). This aspect focuses on the LLM’s capacity for inquiry and exploration within scientific contexts, such as analyzing scientific texts, identifying key concepts, and questioning relevant information.
  • L3: Thinking profoundly (i.e., knowledge reasoning). This criterion examines the model’s capacity for critical thinking, logical deduction, numerical calculation, function prediction, and the ability to engage in reflective reasoning to solve problems.
  • L4: Discerning clearly (i.e., knowledge discernment). This aspect evaluates the LLM’s ability to make correct, secure, and ethical decisions based on scientific knowledge, including assessing the harmfulness and toxicity of information, and understanding the ethical implications and safety concerns related to scientific endeavors.
  • L5: Practicing assiduously (i.e., knowledge application). The final dimension assesses the LLM’s capability to apply scientific knowledge effectively in real-world scenarios, such as solving complex scientific problems and creating innovative solutions.
Building upon the above design philosophy, we develop the SciKnowEval benchmark specifically tailored for assessing multi-level scientific knowledge in LLMs. In particular, we undertake meticulous designs in terms of data scale, diversity and quality when constructing the evaluation dataset:
@@ 3.2 Data Collection Methods
I. Generating New QAs from Literature Corpus.
Literature and textbooks cover a broad range of scientific knowledge, and leveraging this data will facilitate a comprehensive evaluation of LLMs’ capabilities in the scientific domains. We collect massive papers from article preprint platforms (e.g., BioRxiv), literature databases (e.g., PubMed), and textbook databases (e.g., LibreTexts). We utilize LLMs to automate the procedures of QA pair generation. Specifically, following domain experts’ advice, we carefully design effective prompts for literature QA tasks. These prompts exhibited in Appendix A5 guide the LLM to extract relevant professional knowledge from literature and textbook paragraphs, enabling it to generate new QA pairs around this expertise. To ensure quality assessment of the generated questions, we emphasize in the prompts that answers must be explicitly found in the original text without introducing any external information.
II. Refactoring the Existing QAs.
We sample additional QAs from existing open-source scientific benchmarks, including MedMCQA [30], SciEval [34], MMLU [16], XieZhi [13], PubMedQA [17], and HarmfulQA [3]. To mitigate the risk of data contamination and leakage in these benchmarks, we employ LLMs to refactor these QAs in various forms, such as question rewriting and option reordering. Moreover, in cases where some QAs lack explicit annotations indicating their corresponding levels in SciKnowEval, LLMs are utilized to automatically categorize the data into distinct levels.
III. Transforming the Scientific Databases.
To enhance the variety and scope of tasks in our dataset, we select several structured databases and transform them into textual formats suitable for evaluating LLMs. These databases mainly include molecular (e.g., PubChem [19]), protein (e.g., UniProtKB [6]), and cellular-related (e.g., SHARE-seq [26]) sequence information, which contain annotations related to structure, properties, and functions. We can utilize these annotations to construct QA pairs. Specifically, we first conduct preliminary quality screening, such as filtering out chemically invalid SMILES from PubChem using the RDKit library. We then design multiple question templates to transform the structured sequence-annotation pairs into natural language formats, including multiple-choice questions, true/false questions, and short-answer questions.
@@ 3.3 Data Quality Control
The primary concern of data quality is the QAs generated from literature corpus. To ensure the generated data with high quality, we employed a three-stage data screening process:
@@
  • Initial screening by LLMs. We first explicitly instructed LLM during data generation that the correct options must be clearly identifiable from the provided literature snippets. After data generation, we prompt LLMs to simulate an open-book exam task to determines whether each question’s answer can be found in the corresponding snippet.
@@
  • Human evaluation. We randomly selected approximately 5% of the generated questions and provided them with two domain experts. During the evaluation, we used the instructions in Table A11 to guide the evaluators, asking them to thoroughly assess the data and classify it into binary categories of “Yes” and “No” for quality.
  • Post-screening by LLMs. We employed LLMs to summarize the failure types of the identified low-quality entries, and added them into the prompt to conduct a full dataset quality assessment, discarding similar types of low-quality questions.
More details of data quality control are provided in Appendix A7.
3.4 Final Datasets
Based on the above data collection and screening process, we construct the SciKnowEval dataset, consisting of four subsets for Biology, Chemistry, Physics and Materials, respectively. Table 3.4 shows the overall statistics of the constructed dataset. Table 3 provides an overview of the biological dataset, and Table A8 summarizes the datasets for other domains. Detailed question examples are available in Appendix A6.
Table 2: Statistics of the SciKnowEval dataset.
Statistics Number
Total Questions 70,203
Subjects/Tasks 4/78
L1 Questions 39,264 (55.93%)
L2 Questions 12,896 (18.37%)
L3 Questions 8,368 (11.92%)
L4 Questions 5,257 (7.49%)
L5 Questions 4,418 (6.29%)
Statistics Number
Average question length 50.38
Average option length 6.25
Average answer length 56.60
Multiple-choice Questions 52,770 (75.17%)
Constrained Generation Question 10,722 (15.27%)
True or False Questions 4,723 (6.73%)
Relation Extraction Question 1,988 (2.83%)
Table 3: Overview of the biological dataset in SciKnowEval. Abbr., MCQ: multiple choice questions; T/F: true/false; CLS: classification; RE: relation extraction; GEN: generative task. Data collection methods I, II and III are in Fig. 2. The detailed data source listed in Appendix A4. The overview of other domain datasets (chemistry, physics and materials) is presented in Table A8.
Domain Ability Task Name Task Type Data Source Method #Questions
Bio. L1 Biological Literature QA MCQ Literature Corpus I 14,862
Protein Property Identification MCQ UniProtKB III 1,500
Protein Captioning GEN UniProtKB III 930
L2 Drug-Drug Relation Extraction RE Bohrium II 464
Biomedical Judgment and Interpretation T/F PubMedQA II 947
Compound-Disease Relation Extraction RE Bohrium II 500
Gene-Disease Relation Extraction RE Bohrium II 203
Detailed Understanding MCQ LibreTexts I 828
Text Summary GEN LibreTexts I 1,291
Hypothesis Verification T/F LibreTexts I 618
Explanation MCQ LibreTexts I 648
L3 Solubility Prediction MCQ PEER, DeepSol III 207
$\beta$-lactamase Activity Prediction MCQ PEER, Envision III 203
Fluorescence Prediction MCQ PEER, Sarkisyan’s III 203
GB1 Fitness Prediction MCQ PEER, FLIP III 208
Stability Prediction MCQ PEER, Rocklin’s III 204
Protein-Protein Interaction MCQ STRING, SHS27K, SHS148K III 207
Biological Calculation MCQ MedMCQA, SciEval, MMLU II 60
L4 Biological Harmful QA GEN Website I 297
Proteotoxicity Prediction MCQ, T/F UniProtKB III 510
Biological Laboratory Safety Test MCQ, T/F LabExam (ZJU) II 192
L5 Biological Protocol Procedure Design GEN Protocol Journal I 577
Biological Protocol Reagent Design GEN Protocol Journal I 588
Protein Design GEN UniProtKB III 949
Single Cell Analysis GEN SHARE-seq III 300

### v4

# SciKnowEval: A Comprehensive Dataset for Evaluating Scientific Knowledge of Large Language Models

@@ Abstract
Large language models (LLMs) are playing an increasingly important role in scientific research, yet there remains a lack of comprehensive benchmarks to evaluate the breadth and depth of scientific knowledge embedded in these models. To address this gap, we introduce SciKnowEval, a large-scale dataset designed to systematically assess LLMs across five progressive levels of scientific understanding: memory, comprehension, reasoning, discernment, and application. SciKnowEval comprises 28K multi-level questions and solutions spanning biology, chemistry, physics, and materials science. Using this benchmark, we evaluate 20 leading open-source and proprietary LLMs. The results show that while proprietary models often achieve state-of-the-art performance, substantial challenges remain—particularly in scientific reasoning and real-world application. We envision SciKnowEval as a standard benchmark for evaluating scientific capabilities in LLMs and as a catalyst for advancing more capable and reliable scientific language models.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have demonstrated an impressive capability in storing and recalling world knowledge, continuously expanding the boundaries of artificial intelligence. Their exceptional performance has permeated diverse specialized domains, including the scientific domain, leading to the emergence of scientific LLMs, such as Galactica (taylor2022galactica, ), SciGLM (zhang2024sciglm, ), and ChemLLM (zhang2024chemllm, ). To steadily advance scientific research, it is crucial to establish reliable benchmarks that comprehensively evaluate these models’ capability in handling scientific knowledge.
@@
While several existing LLM benchmarks (li2023cmmlu, ; zhong2023agieval, ; clark2018think, ) have incorporated scientific questions into their evaluations, and some benchmarks (sun2024scieval, ; wang2023scibench, ; cai2024sciassess, ; welbl2017crowdsourcing, ; lu2022learn, ; guo2023can, ) are specifically tailored for the scientific domain, we argue that the current benchmarks do not fully evaluate the potential of LLMs in scientific research due to their inherent limitations. Firstly, many existing benchmarks, such as AGIEval (zhong2023agieval, ), SciQ (welbl2017crowdsourcing, ), and ScienceQA (lu2022learn, ), include science questions only up to the high school level, failing to tap into the deeper capability of LLMs. Secondly, recent scientific domain benchmarks like ChemLLMBench (guo2023can, ), SciBench (wang2023scibench, ), and SciAssess (cai2024sciassess, ), despite involving more specialized scientific tasks, lack a comprehensive evaluation system, resulting in a limited understanding of capabilities. Lastly, most benchmarks overlook the assessment of safety issues in scientific research, even those attempting a multi-dimensional comprehensive evaluation such as SciEval (sun2024scieval, ).
Figure 1: Illustration of SciKnowEval. (a) Scientific Domains: Our dataset contains the four subsets of biology, chemistry, material, and physics. (b) Data Sources: We collect our data from various sources, including articles, textbooks, and other sources. (c) Question Types: Our dataset has four types of questions, including relation-extraction questions, multiple-choice questions, content generation, and true/false questions. (d) Five Progressive Levels and Corresponding Examples: We evaluate the LLMs in five ability levels, including their abilities of knowledge memory, comprehension, reasoning, discernment, and application. (e) Question Distribution: The distribution of questions across domains and ability levels.
@@
In response to these deficiencies, we propose a comprehensive Scientific Knowledge Evaluation dataset, referred to as SciKnowEval, as illustrated in Fig. 1. This dataset is designed to assess LLMs based on their proficiency across five progressive levels, with each level offers a unique perspective on evaluating the capabilities of LLMs in handling scientific knowledge, including memory, comprehension, reasoning, discernment, and application. In comparison to existing benchmark datasets, SciKnowEval mainly has the following characteristics: (1) It designs a systematic scientific knowledge evaluation framework that encompasses five progressive levels to mirror the learning process of humans. (2) It uses data from diverse sources, including scientific textbooks, literature, and databases, making it diverse and large-scale. (3) It places significant emphasis on scientific ethics and safety while comprehensively evaluating capabilities.
@@
SciKnowEval represents a comprehensive dataset for assessing the capability of LLMs in processing and utilizing scientific knowledge. It aims to promote the development of scientific LLMs that not only possess extensive knowledge but also demonstrate ethical discernment and practical applicability. The contributions of this paper can be summarized as follows:
@@
  • We construct a large-scale evaluation dataset comprised of 28K diverse scientific problems from the domains of biology, chemistry, physics, and material science, accompanied by corresponding solutions and evaluation metrics, facilitating an extensive assessment of the breadth and depth of scientific knowledge encapsulated in LLMs.
  • We evaluate a wide range of advanced LLMs (including 7 proprietary LLMs, 8 open-source general-purpose LLMs, and 5 scientific LLMs) and rank their performance with the SciKnowEval dataset, elucidating both their strength and weaknesses.
2 Methods
@@ 2.1 Design Philosophy
The profound principles of Confucius inspire the design philosophy of SciKnowEval elucidated in the ancient Chinese book “Doctrine of the Mean” enwiki1 : Studying extensively, Enquiring earnestly, Thinking profoundly, Discerning clearly, and Practicing assiduously. This principle reflects the five progressive levels in the human learning process. Specifically, each level provides a perspective to assess the proficiency of LLMs, as described below.
  • L1: Knowledge Memory. This dimension evaluates an LLM’s ability to store and retrieve a vast range of factual scientific knowledge across multiple domains. It measures the breadth and accuracy of the model’s memory, including definitions, taxonomies, historical facts, and widely accepted scientific principles.
  • L2: Knowledge Comprehension. This aspect focuses on the LLM’s capacity for inquiry and exploration within scientific contexts, such as analyzing scientific texts, identifying key concepts, and questioning relevant information.
  • L3: Knowledge Reasoning. This criterion examines the model’s capacity for critical thinking, logical deduction, numerical calculation, function prediction, and the ability to engage in reflective reasoning to solve problems.
  • L4: Knowledge Discernment. This aspect evaluates the LLM’s ability to make correct, secure, and ethical decisions based on scientific knowledge, including assessing the harmfulness and toxicity of information, and understanding the ethical implications and safety concerns related to scientific endeavors.
  • L5: Knowledge Application. The final dimension assesses the LLM’s capability to apply scientific knowledge effectively in real-world scenarios, such as solving complex scientific problems and creating innovative solutions.
Building upon the above design philosophy, we develop the SciKnowEval dataset specifically tailored for assessing multi-level scientific knowledge in LLMs. In particular, we undertake meticulous designs in terms of data scale, diversity and quality when constructing the evaluation dataset:
@@ 2.2 Data Collection Methods
I. Generating New QAs from Literature Corpus
Literature and textbooks cover a broad range of scientific knowledge, and leveraging this data will facilitate a comprehensive evaluation of LLMs’ capabilities in the scientific domains. We collect massive papers from article preprint platforms (e.g., BioRxiv), literature databases (e.g., PubMed), and textbook databases (e.g., LibreTexts). We utilize LLMs to automate the procedures of QA pair generation. Specifically, following domain experts’ advice, we carefully design effective prompts for literature QA tasks. These prompts exhibited in A5 guide the LLM to extract relevant professional knowledge from literature and textbook paragraphs, enabling it to generate new QA pairs around this expertise. To ensure quality assessment of the generated questions, we emphasize in the prompts that answers must be explicitly found in the original text without introducing any external information.
II. Refactoring the Existing QAs
We sample additional QAs from existing open-source scientific benchmarks, including MedMCQA pal2022medmcqa , SciEval sun2024scieval , MMLU hendrycks2020measuring , XieZhi gu2024xiezhi , PubMedQA jin2019pubmedqa , and HarmfulQA bhardwaj2023red . To mitigate the risk of data contamination and leakage in these benchmarks, we employ LLMs to refactor these QAs in various forms, such as question rewriting and option reordering. Moreover, in cases where some QAs lack explicit annotations indicating their corresponding levels in SciKnowEval, LLMs are utilized to automatically categorize the data into distinct levels.
III. Transforming the Scientific Databases
To enhance the variety and scope of tasks in our dataset, we select several structured databases and transform them into textual formats suitable for evaluating LLMs. These databases mainly include molecular (e.g., PubChem kim2021pubchem ), protein (e.g., UniProtKB uniprot2023uniprot ), and cellular-related (e.g., SHARE-seq ma2020chromatin ) sequence information, which contain annotations related to structure, properties, and functions. We can utilize these annotations to construct QA pairs. Specifically, we first conduct preliminary quality screening, such as filtering out chemically invalid SMILES from PubChem using the RDKit library. We then design multiple question templates to transform the structured sequence-annotation pairs into natural language formats, including multiple-choice questions, true/false questions, and short-answer questions.
@@ 2.3 Data Quality Control
To ensure the generated data with high quality, we employed a three-stage data screening process:
Initial screening by LLMs
@@
Our primary concern is the "Multiple Choice Questions (MCQ)" tasks entirely generated by LLMs, such as the Literature QA task. To ensure the correctness of LLM-generated answers, we first explicitly instructed LLM during data generation that the correct options must be clearly identifiable from the provided literature snippets. After data generation, we prompt GPT-4o (Table A9) to simulate an open-book exam task, where it determines whether each question’s answer could be found in the corresponding literature snippet, and if so, GPT-4o could provide answers to questions based on text snippets (for example, identifying the correct option for multiple-choice questions from a snippet). By comparing these answers with the previously generated answers, we can verify the accuracy of the original answers. Otherwise, we consider the answers to the questions unverifiable, and we simply delete them.
Human evaluation
@@
We randomly selected approximately 5% of the questions from each task and provided them with two experts in biology and chemistry, with the assistance of five graduate students from related fields. It took a week to complete the quality evaluation. During the evaluation, we used the instructions in Table A10 to guide the evaluators, asking them to thoroughly assess the data and classify it into binary categories of "Yes" and "No" for quality. Only data that fully met the requirements was rated "Yes." Ultimately, we identified 2.1% instances of data that were rated "No" after the first human evaluation stage.
Post-screening by LLMs
We employed LLMs to summarize the failure types of these low-quality entries, and added them into the prompt to conduct a full dataset quality assessment, discarding similar types of low-quality questions. We repeated the stage2&3 twice and additionally performed stage 2 one more time. Finally, the low-quality entries identified by experts in stage 2 is less than 0.2%. Since we performed stage 2 three times, each time sampling 5% of the data without replacement from each task, the total amount of data verified for quality exceeded 10% in the end. By implementing these stages, we ensure that the SciKnowEval dataset maintains a high standard of data quality.
2.4 Overview of the SciKnowEval Dataset
The SciKnowEval dataset is constructed by generating new QAs, refactoring existing QAs, and transforming the scientific databases (see Method Section for more details). The dataset consists of four subsets for Biology (28.44%, inclusing 8,076 questions), Chemistry (31.68%, including 8994 questions), Physics (17.5%, including 4,967 questions), and Materials (22.38%, including 6,355 questions), with the task format of multiple-choice questions(65.67%, including 18,670 questions), relation extraction questions(6.12%, including 1,737 questions), true or false questions(11.36%, including 3,228 questions), and generation questions(16.75%, including 4,757 questions). In total, our dataset comprises 58 tasks and 28,392 samples, providing a comprehensive benchmark for evaluating scientific knowledge in LLMs. Table A1 summarizes the datasets for the four domains.